Tickets for a play at the community theater cost $16 for an adult and $2 for a child. If 210 tickets were sold and the receipts were $2380, how many of each type of ticket were sold?
Answer by rmromero(383)


Tickets for a play at the community theater cost $16 for an adult and $2 for a child. If 210 tickets were sold and the receipts were $2380, how many of each type of ticket were sold?

What is asked in the problem?
How many of each type of tickets were sold?
Given:
Tickets for a play at the community theater cost $16 for an adult
and $2 for a child.
If 210 tickets were sold and the receipts were $2380.
Representation:
Let x be the number of tickets sold for adults
y be the number of tickets sold for kids
Equations:
The tickets sold were 210. adult's tickets plus children's ticket
x + y = 210
The receipt cost 2380. Adults' ticket cost $16 and children's $2.
16x + 2y = 2380
Solve the system of equations
x + y = 210
16x + 2y = 2380
___________________

Multiply -2 to x + y = 210 to eliminate y then add the 2 equations
-2x - 2y = -420
16x + 2y = 2380
___________________
14x = 1960 Divide both sides by 14
x = 140 there are 140 adult tickets sold

Choose either of the 2 equation. Substitute the value of x to find y.
x + y = 210
140 + y = 210
y = 210 - 140
y = 70 There are 70 children tickets sold.

Checking
x + y = 210 ; x = 140 & y = 70
140 + 70 = 210
210 = 210 True

16x + 2y = 2380
16 (140) + 2 (70) = 2380
2240 + 140 = 2380
2380 = 2380 True

Therefore the solution we have is correct x = 140 and y = 70. There are 140 adult tickets sold and 70 children tickets sold.

Rmromero